# Anole: Adapting Diverse Compressed Models for Cross-scene Prediction on Mobile Devices

### v1

@@ IV-C Training Decision Model
Specifically, we use $\mathcal{M}_{scene}$ as a backbone neural network to obtain scene representation, denoted as $h^{s}_{i}$, for every data sample $x_{i}\in\Psi_{i}^{sub},i\in[1,n]$. In this way, $h^{s}_{i}$ will retain the scene-related information. The model decision here can be formulated as a multi-class classification problem. The label of $x$ for decision model training is a vector, referred to as a model allocation vector $v^{x}=\{v_{i}^{x},i\in[1,n]\}$, where the $i$-th element $v_{i}^{x}$, denotes whether $x\in\Psi_{i}^{sub}$. The cross entropy loss function [20] is used for training the decision model. Note that during the training of decision model $\mathcal{M}_{decision}$, the parameter of $\mathcal{M}_{scene}$ is frozen to improve training efficiency and enhance the generalization of $\mathcal{M}_{decision}$ [21].
V Online Model Inference
V-A Model Selection Strategy
@@
V-B Cache-based Model Deployment
@@
Given a limited video memory budget, it is tricky to pre-load best models in memory. We examine the utility of 19 YOLOv3-tiny compressed models obtained according to the algorithm stated in §$\S$§IV-A (see §$\S$§VI-A2 for more details) when conducting object detection on the five driving video clips. Figure 4(b) depicts the ratio of being the top one model over all clips for all compressed models. It can be seen that the probability of being the best model follows a power-law distribution. This observation suggests that high-level inference performance can be sustained by deploying only a small number of supreme models. Inspired by this observation, we adopt a Least Frequently Used (LFU) strategy [22] to update models in GPU memory. In the occasion of a model miss, the model with the highest suitability probability in GPU memroy will be used for inference.
VI Evaluation
VI-A Methodology
VI-A1 Datasets
We evaluate Anole on a typical mobile inference task, i.e., vehicle detection on driving videos (VD), based on the following datasets and real-world experiments.
(a) CDF of image brightness
(b) CDF of image contrast
(c) CDF of the number of objects
(d) CDF of the ratio of object area
  • KITTI [23]: comprises 389 stereo and optical flow image pairs, stereo visual odometry sequences of 39.2 km length, and more than 200k 3D object annotations captured in cluttered scenarios (up to 15 cars and 30 pedestrians are visible per image). For online object detection, KITTI consists of 21 training sequences and 29 test sequences.
  • BDD100k [24]: contains over 100k video clips regarding ten autonomous driving tasks. Clips of 720p and 30fps were collected from more than 50 thousand rides in New York city and San Francisco Bay Area, USA. Each clip lasts for 40 seconds and is associated with semantic attributes such as the scene type (e.g., city, streets, residential areas, and highways), weather condition and the time of the day.
  • SHD: contains 100 driving video clips of one minute recorded in March 2022 with a 1080p dashcam in Shanghai city, China. Clips were collected from ten typical scenarios, including highway, typical surface roads, and tunnels, at different time in the day. LabelImg [25] is employed to label objects in all images.
VI-A2 Implementation
@@
Platform CPU GPU GPU Memory Flash/Disk
@@
Jetson TX2 NX ARM A57 Pascal 4GB 32GB
Laptop i7-10750H RTX 2070 8GB 1TB
We implement the offline scene profiling on a server equipped with 128GB RAM and 4 Nvidia 2080 Ti GPUs, running a Linux distribution. We implement online model inference on three typical mobile devices, i.e., a Nvidia Jetson Nano, a Nvidia Jetson TX2 NX and a Windows laptop. Pytorch is employed as the inference engine and TensorRT [26] is used for the run-time acceleration on both Jetson devices, running a Linux distribution. OpenCV is compiled on CPU for balancing the usage of CPU and GPU. The hardware configurations are shown in Table I. ResNet18 [16] and a MLP of two layers are used to train the $\mathcal{M}_{scene}$ and the $\mathcal{M}_{decision}$, respectively. Compressed models for object detection are fine-tuned on YOLOv3-tiny [17] pre-trained on the COCO [27] dataset. Details of all deployed models are listed in Table II. Compressed models are trained with Algorithm 1. A total of 19 compressed models are trained to provide compressed models for inference in all possible scenes.
TABLE II: Details of deployed models, where FLOPS of the deep model YOLOv3 is 10×10$\times$10 × bigger than YOLOv3-tiny and Resnet18.
Model Role FLOPS Weights
YOLOv3-tiny Compress model 5.56 Bn 34 MB
@@
(a) $\mathcal{M}_{scene}$
(b) $\mathcal{M}_{decision}$
@@
VI-A3 Candidate Methods
We compare Anole with the following candidate methods:
  • Single Deep Model (SDM) [17]: One single deep model is trained with all training samples for online inference, i.e., a fully-fledged YOLOv3 is trained.
  • Single Shallow Model (SSM) [28]: One single compressed model is trained with all training samples for online inference, i.e., a YOLOv3-tiny is trained.
  • Clustering-based Domain Generalization (CDG) [29]: Compressed models are trained on domains defined by clustering training data samples in the feature space. During online prediction, the compressed model trained on the cluster which has the closest mean compared with the feature of the test sample is selected for use.
  • Dataset-based Multiple Models (DMM): One separate compressed model is trained on each training dataset, i.e., the KITTI, BDD100k, and SHD datasets. During online prediction, the compressed model corresponding to the same dataset as the test sample is selected for use.
VI-A4 Metrics
[…]
VI-B Effect of Scene Profiling Models
VI-B1 Scene Encoder $\mathcal{M}_{scene}$
We first test $\mathcal{M}_{scene}$ on classifying scenes on the validation set of seen scenes. Scenes are defined based on the multi-level clustering results. Figure 6(a) shows the scene classification confusion matrix of scene encoder $\mathcal{M}_{scene}$ on the validation set. It can be seen that $\mathcal{M}_{scene}$ works well among almost all scenes. There also exist some exceptional scenes that are confusing to $\mathcal{M}_{scene}$. We merge similar scenes in the feature space before training compressed models.
VI-B2 Decision Model $\mathcal{M}_{decision}$
We evaluate the ability of $\mathcal{M}_{decision}$ in selecting the top-one model on the validation set of seen data. Figure 6(b) show the confusion matrix of the $\mathcal{M}_{decision}$ models predicting best models versus true best models. It can be seen that $\mathcal{M}_{decision}$ have basic model selection ability. This is because the decision of model selection is based on the well-trained $\mathcal{M}_{scene}$, with one scene corresponding to a group of suitable models. We can also see that $\mathcal{M}_{decision}$ may make mistakes on some models,. This is because the top one model may not be significantly better than other models.
@@ VI-B2 Decision Model $\mathcal{M}_{decision}$
VI-C Effect of Cache-based Model Update Strategy
To effectively evaluate the effect of our cache-based model update strategy, we synthesize six fast-changing video clips, denoted as T1-T6. Specifically, for each synthesized video clip, we randomly select 5 clips from the 64 clips in the dataset. For each selected clip, we randomly cut a video segment of 100 frames (from the testing set for a seen clip) and then splice all video segments, resulting a synthesized video clip of 500 frames. We then conduct model inference using Anole on T1-T6.
VI-C1 Scene Duration
VI-C2 Cache Miss Rate
VI-D Cross-scene Experiments
In this experiment, we investigate the performance of all candidate methods cross fast-changing scenes, using samples in the test set of seen data. To show the instantaneous performance changes, F1 score is calculated every ten frames.
VI-D1 Performance Comparison
Figure 8 plots the CDFs of F1 score of all candidate methods on each test set of seen data selected from KITTI, BDD100k and SHD, respectively. For both tasks, Anole outwits other methods in terms of accuracy. Moreover, other methods exhibit inconsistent performance across different datasets. For example, DMM gains good performance on the KITTI and SHD datasets, while SDM only performs well on the BDD100K dataset. This discrepancy arises because DMM fits simpler datasets whereas SDM is biased towards BDD100k due to the overwhelming number of training samples.
VI-D2 Effect of Data Segmentation and Model Adaptation
It is a common practice to train an individual model on each datasets (i.e., DMM) or to segment a dataset according to feature similarity and train respective models (i.e., CDG). It can be seen that DMM performs similarly to Anole for simple training datasets such as the KITTI and the SHD datasets, but DMM performs poorly on large and complex datasets like BDD100k. In contrast, CDG trains and selects models on similar data samples. However, the inference accuracy of CDG is not as good as that of Anole over all test sets for both tasks. This demonstrates Proposition 1, which states that a model trained on a scene may not always perform well on that scene. In contrast, Anole employs a decision model to learn the appropriate scenes and determines which model is most suitable for online prediction, resulting in stable performance. Furthermore, although deep-model-based method SDM is generally assumed to have better performance, we surprisingly find that Anole outwits SDM on all test sets. This implies that training a single large DNN model for cross-scene inference is more difficult than training and choosing from a set of specialized compressed models.
VI-E New-scene Experiments
TABLE III: Inference accuracy of all candidate methods obtained on unseen data. The best results are indicated in bold while the second-best results are marked in blue.
Method KITTI BDD100k SHD Mean
St., Da. Ur., Ni. Ur., Da. Hi., Du. St., Ni. Tu., Ni.
@@ VI-E New-scene Experiments
    SDM uses a deep model, resulting in higher latency, larger memory usage (Table IV), and higher power consumption (Figure 11).
(a) Implementation of Anole on Jetson TX2 NX deployed on a UAV.
(b) Results in a typical night scenario.
Figure 9: (a) Implementation of Anole on a Jetson TX2 NX connected with a 1080p HD camera; (b) Visualization of vehicles detected in a night scenario.
In this experiment, we examine the performance of all candidate methods in new scenes, using unseen data. Particularly, six unseen video clips include one clip from KITTI with attributes of {Street, Day}, four scenes from BDD100k with attributes of {Urban, Night}, {Urban, Day}, {Highway, Dusk}, and {Street, Night}, and one scene from SHD with attributes of {Tunnel, Night}. Table III list the accuracy results. It can be seen that though SDM with a much larger model size is expected to excel other shallow-model-based methods on unseen scenes, Anole demonstrates supreme generalization ability and even outperforms SDM on all unseen data. As for unseen scenes from BDD100k, Anole can still achieve high accuracy comparable to that of SDM.
VI-F Real-world Experiments

### v2

@@ V-C Training Decision Model
Specifically, we use $\mathcal{M}_{scene}$ as a backbone neural network to obtain scene representation, denoted as $h^{s}_{i}$, for every data sample $x_{i}\in\Psi_{i}^{sub},i\in[1,n]$. In this way, $h^{s}_{i}$ will retain the scene-related information. The model decision here can be formulated as a multi-class classification problem. The label of $x$ for decision model training is a vector, referred to as a model allocation vector $v^{x}=\{v_{i}^{x},i\in[1,n]\}$, where the $i$-th element $v_{i}^{x}$, denotes whether $x\in\Psi_{i}^{sub}$. The cross entropy loss function [de2005tutorial] is used for training the decision model. Note that during the training of decision model $\mathcal{M}_{decision}$, the parameter of $\mathcal{M}_{scene}$ is frozen to improve training efficiency and enhance the generalization of $\mathcal{M}_{decision}$ [Li2022].
VI Online Model Inference
VI-A Model Selection Strategy
@@
VI-B Cache-based Model Deployment
@@
[…]
VII Evaluation
VII-A Methodology
VII-A1 Tasks and Datasets
We evaluate Anole on two mobile inference tasks, i.e., handwritten digit recognition (HDR), coke grades prediction on mobile agents (CGP) and vehicle detection on driving videos (VD), based on the following datasets and real-world experiments.
  • Handwritten Digit Datasets: The MNIST dataset [lecun1998gradient] contains 70,000 black and white images of handwritten digits. The images are normalized to a $\times28$ pixel bounding box and anti-aliased. As illustrated in Figure 7, we also construct two datasets, i.e., rotated MNIST and noisy MNIST, by applying random rotation ranging from $45^{\circ}$ to $45^{\circ}$ and random gaussian noise with zero mean and a variance of 1 to the original MNIST images, respectively. We mix all three datasets and partition all images into seen and unseen categories with a ratio of 3:2. We further divide seen images into training, validation, and testing sets with a ratio of 2:1:1.
  • Coking Production Decision Making Dataset: The Coke Production Dataset (CPD) is collected from a leading international cloud service provider. It is to predict coke grades for coke production operations where decisions with the highest grades will be taken. A total of 200 coke production experiments collected from July, 2020 to March, 2021 are collected in the CPD dataset. Each sample includes structured features such as the ratio of different coal raw materials and the quality of the refined coke is labeled for prediction. Such a prediction can be used to optimize the decision of industrial production. The samples in the CPD dataset vary from multiple factors such as raw material loactions and production specifications owing to the mobility among production in different batches. We divide the samples into training, validation and testing sets with a ratio of 3:1:1.
[…]
Figure 7: Example images in the original and two constructed MNIST datasets.
@@ VII-A1 Tasks and Datasets
Platform CPU GPU GPU Mem. Flash/Disk
@@
Jetson TX2 ARM A57 Pascal 4GB 32GB
Laptop i7-10750H RTX 2070 8GB 1T
VII-A2 Implementation
We implement the offline scene profiling on a server equipped with 128GB RAM and 4 Nvidia 2080 Ti GPUs, running a Linux distribution. We implement online model inference on three typical mobile devices, i.e., a Nvidia Jetson Nano, a Nvidia Jetson TX2 NX and a Windows laptop. Pytorch is employed as the inference engine and TensorRT [vanholder2016efficient] is used for the run-time acceleration on both Jetson devices, running a Linux distribution. OpenCV is compiled on CPU for balancing the usage of CPU and GPU. The hardware configurations are shown in Table I. For the HDR task, because of small image size, we direct conduct multi-level clustering on image pixels without training the $\mathcal{M}_{scene}$ and use a multi-layer perceptron (MLP) [kruse2022multi] with two layers to train the $\mathcal{M}_{decision}$ for its light-weight characteristic. Compressed models for image classification are trained based on LeNet [lecun1998gradient]. For the CGP task, we directly use the structured features without training the $\mathcal{M}_{scene}$ and use a decision tree (DT) model [breiman2017classification] based on Scikit-Learn [scikit-learn] as $\mathcal{M}_{decision}$ for its superior performance and interpretability. Compressed models for coke grade prediction are trained based on Linear Regression (LinearR) [rifkin2007notes]. For the VD task, ResNet18 [he2016deep] and a MLP of two layers are used to train the $\mathcal{M}_{scene}$ and the $\mathcal{M}_{decision}$, respectively. Compressed models for object detection are fine-tuned on YOLOv3-tiny [adarsh2020yolo] pre-trained on the COCO [lin2014microsoft] dataset. Details of all deployed models are listed in Table II. Compressed models of HDR and VD tasks are trained with Algorithm 1 consodering the heavy training overhead of DNN model and compressed models of CGD tasks are trained with Algorithm 2 for a better performance. A total of 25, 17 and 19 compressed models are trained for the HDR, CGD and VD tasks, respectively, to provide compressed models for inference in all possible scenes.
TABLE II: Details of deployed models, where FLOPS of the deep model YOLOv3 is $10\times$ bigger than YOLOv3-tiny and Resnet18.
Task Model Role FLOPS Weights
HDR LeNet Compress model 31 M 234 KB
MLP $\mathcal{M}_{decision}$ 3.6 M 935 KB
Resnet18 Deep model 4.69 Bn 44 MB
Resnet50 Deep model 9.74 Bn 87 MB
CGP LinearR Compress model 0.2 K 0.1 KB
DecisionTree $\mathcal{M}_{decision}$ 0.1 K 1.5 KB
VD YOLOv3-tiny Compress model 5.56 Bn 34 MB
@@ VII-A2 Implementation
VII-A3 Candidate Methods
We compare Anole with the following candidate methods:
  • Single Deep Model (SDM) [adarsh2020yolo]: One single deep model is trained with all training samples for online inference. For the HDR task, a ResNet18 and a ResNet50 are respectively trained for their supreme image classification performance. For the VD task, a fully-fledged YOLOv3 is trained. For the CGP task, no deep models are considered because there is no large model good enough so far for this task.
  • Single Shallow Model (SSM) [redmon2018yolov3]: One single compressed model is trained with all training samples for online inference. For the HDR task, a LeNet is trained. For the VD task, a YOLOv3-tiny is trained.
  • Clustering-based Domain Generalization (CDG) [zheng2019metadata]: Compressed models are trained on domains defined by clustering training data samples in the feature space. During online prediction, the compressed model trained on the cluster which has the closest mean compared with the feature of the test sample is selected for use.
  • Dataset-based Multiple Models (DMM): One separate compressed model is trained on each training dataset, i.e., the original, rotated, and noisy MNIST for the HDR task, the KITTI, BDD100k, and SHD for the VD task, the human-defined subsets for CGP task. During online prediction, the compressed model corresponding to the same dataset as the test sample is selected for use.
VII-A4 Metrics
We evaluate the performance of all candidate methods with respect to inference accuracy and latency. For the HDR task, accuracy is defined as the proportion of correctly predicted samples to the total number of test samples. For the CGD task, we use Error Rate (ER), defined as $\text{ER}=\frac{\sum_{n=1}^{N}|\hat{y}_{n}-y_{n}|}{\sum_{n=1}^{N}y_{n}}$, Mean Absolute Error (MAE), defined as $\text{MAE}=\frac{\sum_{i=1}^{N}|\hat{y}_{i}-y_{i}|}{N}$, Mean Squared Error (MSE), defined as $\text{MSE}=\frac{\sum_{i=1}^{N}(\hat{y}_{i}-y_{i})^{2}}{N_{j}}$ and Pearson Correlation Coefficient (PCCS), defined as $\text{PCCS}=\frac{\sum_{i=1}^{N}(\hat{y}_{i}-\hat{\mu})(y_{i}-\mu)}{\sqrt{\sum_{i=1}^{N}(\hat{y}_{i}-\hat{\mu})^{2}}\sqrt{\sum_{i=1}^{N}(y_{i}-\mu)^{2}}}$ for evaluation, where $N$ denotes the number of all testing data; $\hat{y}_{n}$ and $y_{n}$ are the estimation and the ground truth of the $n$-th sample; $\hat{\mu}^{j}$ and $\mu^{j}$ are the mean of $\hat{y}_{n}$ and $y_{n}$, respectively. For the VD task, we use F1 score, defined as $\text{F1}=\frac{2\cdot p\cdot r}{p+r}$, where $p$ and $r$ denote the precision and recall of detection, respectively. We also consider the end-to-end delay, i.e., the time duration from receiving a test sample to obtaining the corresponding inference result.
VII-B Effect of Scene Profiling Models
VII-B1 Scene Encoder $\mathcal{M}_{scene}$
For the VD task, we test $\mathcal{M}_{scene}$ on classifying scenes on the validation set of seen scenes. Scenes are defined based on the multi-level clustering results. Figure 8(a) shows the scene classification confusion matrix of scene encoder $\mathcal{M}_{scene}$ on the validation set. It can be seen that $\mathcal{M}_{scene}$ works well among almost all scenes. There also exist some exceptional scenes that are confusing to $\mathcal{M}_{scene}$. We merge similar scenes in the feature space before training compressed models.
(a) $\mathcal{M}_{scene}$ for VD
(b) $\mathcal{M}_{decision}$ for HDR
(c) $\mathcal{M}_{decision}$ for VD
@@ VII-B1 Scene Encoder $\mathcal{M}_{scene}$
VII-B2 Decision Model $\mathcal{M}_{decision}$
We evaluate the ability of $\mathcal{M}_{decision}$ in selecting the top-one model on the validation set of seen data. Figure 8(b) and Figure 8(c) show the confusion matrix of the $\mathcal{M}_{decision}$ models predicting best models versus true best models for the HDR and VD tasks, respectively. It can be seen that $\mathcal{M}_{decision}$ have basic model selection ability. This is because the decision of model selection is based on the well-trained $\mathcal{M}_{scene}$, with one scene corresponding to a group of suitable models. We can also see that $\mathcal{M}_{decision}$ may make mistakes on some models,. This is because the top one model may not be significantly better than other models.
VII-C Effect of Cache-based Model Update Strategy
To effectively evaluate the effect of our cache-based model update strategy, we synthesize six fast-changing video clips, denoted as T1-T6, for the VD task. Specifically, for each synthesized video clip, we randomly select 5 clips from the 64 clips in the dataset. For each selected clip, we randomly cut a video segment of 100 frames (from the testing set for a seen clip) and then splice all video segments, resulting a synthesized video clip of 500 frames. We then conduct model inference using Anole on T1-T6.
@@ VII-C Effect of Cache-based Model Update Strategy
VII-C1 Scene Duration
VII-C2 Cache Miss Rate
TABLE III: Inference accuracy of candidate methods achieved in cross-scene and new-scene scenarios on the HDR task.
Method Model
GPU Mem.
Cross-scene New-scene
Noisy Rota. Noisy Rota.
SDM Res.18 45.27 0.773 0.828 0.777 0.826
Res.50 98.31 0.704 0.810 0.702 0.808
SSM LeNet 0.18 0.755 0.751 0.754 0.752
CDG LeNet 0.18 0.099 0.721 0.099 0.721
DMM LeNet 0.18 0.620 0.755 0.619 0.758
Anole LeNet 0.18 0.822 0.844 0.821 0.846
TABLE IV: Performance of different methods on the CGD task.
[] Method1 Latency ER MAE MSE PCCS SSM 0.87s 2.28 1.48 2.69 0.17 CDG 0.87s 5.05 3.27 14.16 0.12 DMM 0.90s 2.14 1.38 2.34 0.16 Anole 1.05s 1.38 0.90 1.40 0.40
  • 1
    There is no effective deep model (i.e., SDM) so far for the CGD task owing to the lacking of large-scale datasets.
VII-D Cross-scene Experiments
In this experiment, we investigate the performance of all candidate methods cross fast-changing scenes, using samples in the test set of seen data. For the VD task, F1 score is calculated every ten frames to show the instantaneous performance changes.
VII-D1 Performance Comparison
Table III lists the accuracy of all candidate methods over all cross-scene HDR test images. Table IV list the considered metrics of all candidate methods on the testing samples of CGD task. Figure 10 plots the CDFs of F1 score of all candidate methods on each test set of seen data selected from KITTI, BDD100k and SHD, respectively. For both tasks, Anole outwits other methods in terms of accuracy. For example, on the HDR task, Anole can even outwits SDM by 6.3%; on the CGD task, Anole can achieve a prediction PCCS increase by 135.5%; on the VD task, ANole can outperform SDM in terms of SDM by over 15%. Moreover, other methods exhibit inconsistent performance across different datasets. For example of the VD task, DMM gains good performance on the KITTI and SHD datasets, while SDM only performs well on the BDD100K dataset. This discrepancy arises because DMM fits simpler datasets whereas SDM is biased towards BDD100k due to the overwhelming number of training samples.
@@ VII-D1 Performance Comparison
VII-D2 Effect of Data Segmentation and Model Adaptation
It is a common practice to train an individual model on each datasets (i.e., DMM) or to segment a dataset according to feature similarity and train respective models (i.e., CDG). It can be seen that DMM performs similarly to Anole for simple training datasets such as all MNIST datasets for the HDR task, and the KITTI and the SHD datasets for the VD task, but DMM performs poorly on large and complex datasets like BDD100k. In contrast, CDG trains and selects models on similar data samples. However, the inference accuracy of CDG is not as good as that of Anole over all test sets for both tasks. This demonstrates Proposition 1, which states that a model trained on a scene may not always perform well on that scene. In contrast, Anole employs a decision model to learn the appropriate scenes and determines which model is most suitable for online prediction, resulting in stable performance. Furthermore, although deep-model-based method SDM is generally assumed to have better performance, we surprisingly find that Anole outwits SDM on all test sets. This implies that training a single large DNN model for cross-scene inference is more difficult than training and choosing from a set of specialized compressed models.
TABLE V: Inference accuracy of all candidate methods obtained on unseen data. The best results are indicated in bold while the second-best results are marked in blue.
Method KITTI BDD100k SHD Average
Street, Day Urban, Night Urban, Day Highway, Dusk Street, Night Tunnel, Night
@@ VII-D2 Effect of Data Segmentation and Model Adaptation
    SDM uses a deep model, resulting in higher latency, larger memory usage (Table VI), and higher power consumption (Figure 13).
VII-E New-scene Experiments
In this experiment, we examine the performance of all candidate methods in new scenes, using unseen data. Particularly, for the VD task, six unseen video clips include one clip from KITTI with attributes of {Street, Day}, four scenes from BDD100k with attributes of {Urban, Night}, {Urban, Day}, {Highway, Dusk}, and {Street, Night}, and one scene from SHD with attributes of {Tunnel, Night}. Table III and Table V list the accuracy results for the HDR and VD tasks, respectively. It can be seen that though SDM with a much larger model size is expected to excel other shallow-model-based methods on unseen scenes, Anole demonstrates supreme generalization ability and even outperforms SDM on all unseen data. As for unseen scenes from BDD100k, Anole can still achieve high accuracy comparable to that of SDM.
VII-F Real-world Experiments
(a) Implementation of Anole on Jetson TX2 NX deployed on a UAV
(b) Results in a typical night scenario
Figure 11: (a) Implementation of Anole on a Jetson TX2 NX connected with a 1080p HD camera. (b) Visualization of vehicles detected in a night scenario.
As depicted in Figure 11(a), we implement all methods on the Nvidia Jetson TX2 NX connected with a 1080p HD camera to conduct real-world experiments in Shanghai city. Well-trained compressed models and the decision model are downloaded to the Jetson device. We conduct real-world experiments in seven driving scenarios with different road conditions and different time in a day. LabelImg [tzutalin2015labelimg] is used to label all recorded frames as the ground truth for offline analysis. Figure 12 plots the F1 score of all methods. Anole outperforms all other candidate methods in all test scenarios. We visualize the car detection results of Anole (white solid frames) and SDM (red dashed frames) in a typical night driving scenario in Figure 11(b). The inference results obtained using SDM frequently contain errors, especially false negative errors as shown in the enlarged subgraph.